OpenStudy (anonymous):

PLEASE HELP QUICKLY! Set up an equation and solve the problem. It takes Terry 2 hours longer to do a certain job than it takes Tom. They worked together for 2 hours; then Tom left and Terry finished the job in 1 hour. How long would it take each of them to do the job alone?

OpenStudy (tkhunny):

Very much first, most important - Write down definitions! Y = Terry's Hours / Job M = Tom's Hours / Job 1/Y = Terry's Job / Hour 1/M = Tome Job / Hour Worked together for 2 hours (2 hr)(1/Y + 1/M) + (1 hr)(1/Y) = 1 Job

OpenStudy (anonymous):

Ok.. i'm a bit confused though. What is the exact answers?

OpenStudy (anonymous):

are*

OpenStudy (tkhunny):

That is the algebra part. You must work through it.

OpenStudy (anonymous):

Terry's hour is y and Tom's hour is x -> y = x + 2 ( hr) The portion both work for 2 hours: = 2/x + 2/(x +2) = .... The remained portion: = 1 - ...

OpenStudy (anonymous):

I have no idea how to solve. I tried and I only have 12 mins to get the right answer.

OpenStudy (anonymous):

You just posted it for almost half an hour, I don't think you have enough time :s

OpenStudy (anonymous):

Can you just give me the answer please?